Question 7 options: A government agency would like to know the average weight of adult catfish in the wild. The need to know how many fish to sample. The population standard deviation is known to be 16.8 lbs from previous studies.

(a) Assuming that they would like to be 90% confident of the results, what is the critical value that corresponds to this given level of confidence? Round to two decimals and remember that critical values are defined to be positive.

(b) If they would like to have a maximum error of 1.4 lbs, how many fish should be sampled?
Answer #1

populaiton sd = 16.8

(a) Critical value for 90% confidence level = 1.645

Critical value = 1.65

(b) E = 1.4

E = z *\frac{ \sigma}{\sqrt{n}}

1.4 = 1.645 *\frac{ 16.8}{\sqrt{n}}

n = 389.6676

n = 390